# PAC Privacy Preserving Diffusion Models

### v1

@@ 4.1 Datasets
Our experiments were carried out using the CelebA dataset (Liu et al., 2015). We specifically targeted attributes that have a balanced distribution of positive and negative samples, as noted in (Rudd et al., 2016), to facilitate the training of classifiers. Consequently, we selected attributes like gender, attractiveness, and smile from the CelebA dataset (referred to as CelebA-gender, CelebA-attractive, and CelebA-smile, respectively) for sampling in Algorithm 1. All the images used in these experiments were of the resolution $64\times 64$.
@@ 4.3 Evaluation Metrics
Table 1: Perceptual score comparisons on CelebA dataset with image resolution $64\times 64$. In our model, we train with data on a specific label respectively, and Model-Gender means our model is trained with CelebA-Gender dataset. The same applies to the following.
Methods $\varepsilon$ FID $\downarrow$ IS $\uparrow$
P3DM-Gender $\varepsilon=1$ 284 1.01
P3DM-Attractive $\varepsilon=1$ 285 1.02
P3DM-Smile $\varepsilon=1$ 284 1.01
DPGEN $\varepsilon=1$ 288 1.01
P3DM-Gender $\varepsilon=10$ 39.3 1.06
P3DM-Attractive $\varepsilon=10$ 44 1.06
P3DM-Smile $\varepsilon=10$ 37.5 1.06
DPGEN $\varepsilon=10$ 37.52 1.04
DPDM $\varepsilon=10$ 118 1.05
(a) DPDM for $\varepsilon=10$
(b) DPGEN for $\varepsilon=10$
(c) P3DM (Ours) for $\varepsilon=10$
@@
Table 2: Privacy Score ($\uparrow$) On Different Models
Dataset $\varepsilon$ P3DM DPGEN DPDM
CelebA-Gender $\varepsilon=1$ 0.45 0.35 -
CelebA-Smile $\varepsilon=1$ 0.425 0.35 -
CelebA-Attractive $\varepsilon=1$ 0.375 0.35 -
CelebA-Gender $\varepsilon=10$ 0.56 0.45 0.225
CelebA-Smile $\varepsilon=10$ 0.37 0.17 0.55
CelebA-Attractive $\varepsilon=10$ 0.3 0.15 0.475
Table 3: $\mathbb{E}||B||_{2}$ to ensure $MI(X;M(X)+B)\leq 1$
@@
P3DM-Gender $\varepsilon=1$ 283.03
P3DM-Attractive $\varepsilon=1$ 279.69
P3DM-Smile $\varepsilon=1$ 280.60
DPGEN $\varepsilon=1$ 281.3
P3DM-Gender $\varepsilon=10$ 328.80
P3DM-Attractive $\varepsilon=10$ 324.44
P3DM-Smile $\varepsilon=10$ 327.98
DPGEN $\varepsilon=10$ 329.48
DPDM $\varepsilon=10$ 315.83
4.4 Empirical Results And Analysis
It is important to note that the $\varepsilon$ presented in our model within the tables below is merely a hyperparameter derived from the Randomized Response (RR) as indicated in Eqn. 3. This $\varepsilon$ should not be confused with the $\varepsilon$ from $\varepsilon$-Differential Privacy (DP), as our model uses PAC privacy rather than $\varepsilon$-DP. We will later illustrate how our model assures privacy through the automatic control of mutual information for a PAC privacy guarantee in this section.
Table 1 details the evaluation results of image visual quality on the CelebA dataset with a resolution of $64\times 64$, while Table 2 shows the privacy scores as calculated by Algorithm 2 for different models. By examining these tables, we can see that our model achieves image quality comparable to the state-of-the-art model DPGEN (Chen et al., 2022), a conclusion also supported by Fig. 2. Additionally, our model registers the highest performance in privacy score, signaling an enhancement in our model’s ability to preserve privacy without substantially impacting image utility.
Furthermore, in Table 3, we compute the multivariate Gaussian matrix $B$ with a $1-\gamma$ noise determination for the deterministic mechanism $M$ followed the work by (Xiao & Devadas, 2023). From this, we show that under the same confidence level of $0.991-\gamma=$ 0.99 to ensure $MI(X;M(X)+B)\leq 1$, the PAC Privacy Preserving Diffusion Model exhibits the smallest norm on $\mathbb{E}||B||_{2}$. This demonstrates that data processed by our model retains the least mutual information with the original dataset, thus affirming our model’s superior performance in privacy protection.
@@ 5 Conclusion
Potential future work includes a deeper exploration of our model’s privacy capabilities. This will involve deploying membership inference attacks, as referenced in works like (Shokri et al., 2017) and (Matsumoto et al., 2023), to assess the resilience of different models. It would also be intereting future work to extend the evaluation of our model’s performance to more complex datasets, with CIFAR-10 (Krizhevsky et al., 2009) and CUB (Wah et al., 2011) being a prime example. These steps are critical for validating and enhancing the robustness of our privacy-preserving model in the face of evolving datasets and attacking approaches.

### v3

@@ 5.1 Datasets
Our experiments were carried out using the CelebA (Liu et al., 2015) datasets. We specifically targeted attributes that have a balanced distribution of positive and negative samples, as noted in Rudd et al. (2016), to facilitate the training of classifiers. Consequently, we selected attributes like gender and smile from the CelebA dataset (referred to as CelebA-gender and CelebA-smile, respectively) for sampling in Algorithm 1. All the images used in these experiments were of the resolution $64\times 64$.
@@ 5.3 Evaluation Metrics
Table 1: Perceptual and privacy score comparisons on CelebA with image resolution $64\times 64$. In our model, we train with data on a specific label respectively, and Model-Gender means our model is trained with CelebA-Gender dataset. The same applies to the following.
P3DM-Gender P3DM-Smile DPGEN DPDM
$\varepsilon$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$
1 175$\pm$2.5 0.6$\pm$0.05 170$\pm$1.85 0.63$\pm$0.08
5 125.8$\pm$3.6 0.55$\pm$0.082 107.85$\pm$4.7 0.6$\pm$0.075 155$\pm$1.5 0.46 $\pm$ 0.076 170$\pm$2 0.57$\pm$0.06
10 44.82$\pm$4.48 0.5$\pm$0.05 37.96$\pm$1.85 0.56$\pm$0.045 39.16$\pm$0.68 0.4$\pm$0.03 117$\pm$1.5 0.47$\pm$0.04
015 40$\pm$0.5 0.45$\pm$0.05 37.9$\pm$0.6 0.5$\pm$0.025 38.5$\pm$0.73 0.4$\pm$0.06 113$\pm$1.2 0.35$\pm$0.05
$\infty$ 34.64$\pm$2.18 0.4$\pm$0.04 36.03$\pm$1.32 0.48$\pm$0.055 34.4$\pm$1.8 0.376$\pm$0.03 110$\pm$1.2 0.3$\pm$0.06
5.4 Empirical Results And Analysis
It is important to note that the $\varepsilon$ presented in our model within the tables below is merely a hyperparameter derived from the Randomized Response (RR) as indicated in Eqn. 4. This $\varepsilon$ should not be confused with the $\varepsilon$ from $\varepsilon$-Differential Privacy (DP), as our model uses PAC privacy rather than $\varepsilon$-DP. We will later illustrate how our model assures privacy through the automatic control of mutual information for a PAC privacy guarantee in this section.
(a) DPDM FID = 117, Privacy Score = 0.47
(b) DPGEN FID = 39.16, Privacy Score = 0.4
(c) P3DM(Ours) FID = 37.96, Privacy Score = 0.56
@@ 5.4 Empirical Results And Analysis
Table 2: $\mathbb{E}||B||_{2}$ to ensure $MI(X;M(X)+B)\leq 1$ on CelebA dataset
@@
P3DM-Gender $5$ 283.03$\pm$2.25
P3DM-Smile $5$ 280.60$\pm$2.57
DPGEN $5$ 281.3$\pm$2.78
P3DM-Gender $10$ 328.80$\pm$1.24
P3DM-Smile $10$ 327.98$\pm$1.45
DPGEN $10$ 329.48$\pm$1.6
DPDM $10$ 335.83$\pm$1.21
P3DM-Gender $\infty$ 332.87$\pm$1.15
P3DM-Smile $\infty$ 331.67$\pm$1.3
DPGEN $\infty$ 332.02$\pm$1.32
DPDM $\infty$ 335.83$\pm$1.56
Table 1 and Fig.3 details the evaluation results of image visual quality and privacy score on the CelebA dataset with a resolution of $64\times 64$, where each datapoint in the figure, or each entry in the table, consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. By examining the table, we can see that our model achieves image quality comparable to the state-of-the-art model DPGEN (Chen et al., 2022), a conclusion also supported by Fig.1.
Additionally, from Fig.3, our model registers the highest performance in privacy score when having similar FIDs, signaling an enhancement in our model’s ability to preserve privacy without substantially impacting image utility. The assertion is further supported by Fig. 2, wherein the CelebA dataset is filtered based on the ’smile’ attribute. Even upon querying the nearest images of P3DM samples, we can distinctly observe that while all P3DM samples exhibit the absence of a smile, the nearest neighbors predominantly display smiling faces. This highlights that, in contrast to other models, the closest images between the generated dataset and the ground truth datasets are notably similar in terms of the ’smile’ expression, while the images generated by our model demonstrate distinctive features. Hence, the P3DM model effectively conceals the ’smile’ attribute.
(a) DPGEN for $\varepsilon=\infty$
(b) DPDM for $\varepsilon=10$
(c) DPGEN for $\varepsilon=10$
(d) P3DM-Smile (Ours) for $\varepsilon=10$
Figure 2: Generated images (the second row) and their nearest neighbors measured by the $l_{2}$ distance between images from CelebA-smile dataset (the first row), with image resolution 64 × 64.
Furthermore, in Table 2, we compute the multivariate Gaussian matrix $B$ with a $1-\gamma$ noise determination for the deterministic mechanism $M$ followed the work by Xiao & Devadas (2023). From this, we show that under the same confidence level of $0.991-\gamma=$ 0.99 to ensure $MI(X;M(X)+B)\leq 1$, the PAC Privacy Preserving Diffusion Model exhibits the smallest norm on $\mathbb{E}||B||_{2}$. This demonstrates that data processed by our model retains the least mutual information with the original dataset, thus affirming our model’s superior performance in privacy protection.
Figure 3: Privacy score and FID curve of all datapoints from different models. Each datapoint in the figure consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. Top-right corner is preferred. The curve from our method, pushes the frontier to the upper-right over DPGEN (Chen et al., 2022) and DPDM (Dockhorn et al., 2022).
6 Limitations
The dataset attributes analyzed in our study predominantly consist of features with balanced positive and negative samples. Attributes characterized by imbalanced distributions remain underexplored. Furthermore, it should be noted that not all datasets include annotated attributes. Therefore, utilizing image-to-text models such as CLIP (Radford et al., 2021) may prove beneficial for the generation of new attribute labels. Lastly, our model faces challenges in synthesizing higher-resolution privacy-preserving images (e.g., 256x256 pixels) and handling complex datasets such as the CUB dataset (Wah et al., 2011)).